OpenStudy (anonymous):

The value of a certain fraction is 3/5. If both the numerator and denominator are increased by 5 , the new fraction is equivalent to 7/10. Find the original fraction.

OpenStudy (anonymous):

hmmm, what's the first step?

OpenStudy (anonymous):

i said 3/5 = N/D and then (N+5) / (D+5) = 7/10 as well

OpenStudy (anonymous):

but if you like, it might be easier just to write (3x) / (5x) as your original, and use (3x+5) / (5x+5) = 7/10

OpenStudy (anonymous):

Thank you!